OpenStudy (anonymous):

Write an exponential function showing the relationship between y and t. A principal of $1500 was invested at 5.25% interest, compounded annually. Let t be the number of years since the start of the investment. Let y be the value of the investment, in dollars

OpenStudy (anonymous):

@Seratul help?

OpenStudy (anonymous):

@rebeccaxhawaii help?

OpenStudy (anonymous):

@phi help?

OpenStudy (phi):

there is a formula for savings growing at some interest rate. do you know it?

OpenStudy (anonymous):

no

OpenStudy (phi):

it is \[ y= P(1+ r)^t\] where P is the starting amount (the principal) and r is the rate (as a decimal

OpenStudy (anonymous):

so would the answer be y = 1500(1 + 0.0525) ?

OpenStudy (anonymous):

y = 1500 (1+ 0.0525)^t

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, but you probably simplify 1+0.0525 to 1.0525 so the answer is \[ y = 1500\cdot 1.0525^t\]

OpenStudy (anonymous):

thank you!